OpenStudy (anonymous):

A square has a diagonal length 13m. What is the side length of the square?

OpenStudy (jdoe0001):

have you covered the 45-45-90 rule?

OpenStudy (anonymous):

a^2 + b^2 = c^2 a^2 + b^2 = 13^2 a = b

OpenStudy (jdoe0001):

keep in mind that both sides are equal, thus a = b so \(\bf c^2 = a^2 + b^2 \implies c^2 = a^2 + a^2 \implies c^2 = 2a^2\) so you have the diagonal, you know what "c" is, then solve for "a"